# Conditional Logical Message Passing Transformer for Complex Query Answering

### v1

@@ Abstract.
Complex Query Answering (CQA) over Knowledge Graphs (KGs) is a challenging task. Given that KGs are usually incomplete, neural models are proposed to solve CQA by performing multi-hop logical reasoning. However, most of them cannot perform well on both one-hop and multi-hop queries simultaneously. Recent work proposes a logical message passing mechanism based on the pre-trained neural link predictors. While effective on both one-hop and multi-hop queries, it ignores the difference between the constant and variable nodes in a query graph. In addition, during the node embedding update stage, this mechanism cannot dynamically measure the importance of different messages, and whether it can capture the implicit logical dependencies related to a node and received messages remains unclear. In this paper, we propose Conditional Logical Message Passing Transformer (CLMPT), which considers the difference between constants and variables in the case of using pre-trained neural link predictors and performs message passing conditionally on the node type. We empirically verified that this approach can reduce computational costs without affecting performance. Furthermore, CLMPT uses the transformer to aggregate received messages and update the corresponding node embedding. Through the self-attention mechanism, CLMPT can assign adaptive weights to elements in an input set consisting of received messages and the corresponding node and explicitly model logical dependencies between various elements. Experimental results show that CLMPT is a new state-of-the-art neural CQA model.
Knowledge Graph, Logical Reasoning, Complex Query Answering, Graph Neural Network
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Reasoning about belief and knowledgeccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Data mining
@@ 1. Introduction
Recently, neural models (Ren and Leskovec, 2020; Amayuelas et al., 2022; Zhang et al., 2021) have been proposed to solve Complex Query Answering (CQA) over incomplete KGs. The complex queries that these models aim to address belong to an important subset of the first-order queries, specifically Existentially quantified First Order queries with a single free variable (EFO-1) (Wang et al., 2022, 2023b). Such EFO-1 queries contain logical operators, including conjunction, disjunction, and negation, as well as the existential quantifier ($\exists$) (Ren and Leskovec, 2020). As shown in Figure 1, one can get the corresponding EFO-1 formula and query graph given a question. The query graph is a graphical representation of the EFO-1 query, where each edge is an atomic formula that contains a predicate with a (possible) negation operator, and each node represents an input constant entity or a variable. Most neural CQA models convert the query graph into the computation graph, where logical operators are replaced with corresponding set operators. These models embed the entity set into specific vector space (Ren and Leskovec, 2020; Zhang et al., 2021; Chen et al., 2022; Yang et al., 2022) and execute set operations parameterized by the neural networks according to the computation graph. However, these models cannot perform well on both one-hop and multi-hop queries simultaneously. They tend to be less effective than classical neural link predictors (Trouillon et al., 2016) on one-hop atomic queries (Arakelyan et al., 2020; Wang et al., 2023b).
@@
To this end, Logical Message Passing Neural Network (LMPNN) (Wang et al., 2023b) proposes a message passing framework based on pre-trained neural link predictors. Specifically, for each edge in the query graph, LMPNN uses the pre-trained neural link predictor to infer an intermediate embedding for a node given neighborhood information. The intermediate embedding can be interpreted as a logical message passed by the neighbor on the edge. Figure 1 demonstrates the logical message passing with red/blue arrows. Then, LMPNN aggregates messages and updates node embeddings in a manner similar to Graph Isomorphism Network (GIN) (Xu et al., 2019). While effective on both one-hop and multi-hop queries, LMPNN ignores the difference between constant and variable nodes. Logical message passing can be viewed as learning an embedding for the variable by utilizing information of constant entities. It makes this embedding similar or close to the embedding of the solution entity in the embedding space of the pre-trained neural link predictor. However, as shown by the red arrows in Figure 1, LMPNN also passes messages to constant entities with pre-trained embeddings and updates their embeddings, which we believe is unnecessary and experimentally demonstrated. In addition, for node embedding updating, despite the expressive power of GIN in graph structure learning (Xu et al., 2019), it cannot dynamically measure the importance of different logical messages. Furthermore, it is unclear whether the GIN-like approach can capture the implicit complex logical dependencies between messages received by a node and between messages and that node. Explicitly modeling the above implicit logical dependencies while being able to measure the importance of different messages dynamically is an open challenge we take on in this paper.
@@
We propose Conditional Logical Message Passing Transformer (CLMPT), a special Graph Neural Network (GNN) for CQA. Each CLMPT layer has two stages: (1) passing the logical message of each node to all of its neighbors that are not constant entity nodes; (2) updating the embedding of each variable node based on the logical messages received by the variable node. This means that during forward message passing, CLMPT does not use the pre-trained neural link predictor to infer the intermediate embeddings (i.e. logical messages) for constant entity nodes with pre-trained embeddings and does not update constant node embeddings. We call such a mechanism Conditional Logical Message Passing, where the condition depends on whether the node type is constant or variable. We empirically verified that it can reduce computational costs without affecting performance. Furthermore, CLMPT uses the transformer (Vaswani et al., 2017) to aggregate messages and update the corresponding node embedding. The self-attention mechanism in the transformer can dynamically measure the importance of different elements in the input and capture the interactions between any two elements. It allows us to explicitly model the logical dependencies between messages received by a node and between messages and that node. We conducted experiments on three popular KGs: FB15k (Bordes et al., 2013), FB15k-237 (Toutanova and Chen, 2015), and NELL995 (Xiong et al., 2017a). The experimental results show that CLMPT is a new state-of-the-art neural CQA model.
@@ 2. Related Work
Complex queries over KGs can be regarded as one-hop atomic queries combined with existential first-order logic operators. The scope of complex queries that existing works can answer is expanded from conjunctive queries (Hamilton et al., 2018; Kotnis et al., 2021), to Existential Positive First-Order (EPFO) queries (Ren* et al., 2020; Arakelyan et al., 2020), and more recently to the existential first-order (Ren and Leskovec, 2020), or more specifically to EFO-1 queries (Wang et al., 2022, 2023b). Recently, neural query embedding (QE) models have been proposed to answer complex queries. Most QE models embed entity sets into specific continuous vector spaces with various forms, including geometric shapes (Choudhary et al., 2021b; Amayuelas et al., 2022; Hamilton et al., 2018; Ren* et al., 2020; Zhang et al., 2021; Liu et al., 2021; Nguyen et al., 2023), probabilistic distributions (Ren and Leskovec, 2020; Yang et al., 2022; Choudhary et al., 2021a), fuzzy logic (Chen et al., 2022), and bounded histograms on uniform grids (Wang et al., 2023a). They represent the complex query in the form of the computation graph and perform neural set operations in the corresponding embedding spaces according to the computation graph. Such neural set operations are usually modeled by complex neural networks (Tolstikhin et al., 2021; Vaswani et al., 2017). In spite of getting query embedding with neural set operations, other QE models represent the queries as graphs to embed queries with GNNs (Daza and Cochez, 2020; Alivanistos et al., 2022) or Transformers (Kotnis et al., 2021; Liu et al., 2022; Wang et al., 2023c). However, these graph representation learning models only focus on EPFO queries, and how they handle the logical negation operator is unclear. Moreover, the aforementioned neural models are still less effective than a simple neural link predictor (Trouillon et al., 2016) on one-hop atomic queries.
@@
There is also a class of neural CQA models based on pre-trained neural link predictors. CQD-CO (Arakelyan et al., 2020) proposes a continuous optimization strategy that uses pre-trained neural link predictors (Trouillon et al., 2016) and logical t-norms (Klement et al., 2013) to solve complex queries. However, according to (Wang et al., 2023b), CQD-CO performs badly on negative queries. LMPNN (Wang et al., 2023b), which is most related to our work, proposes a logical message passing framework based on pre-trained neural link predictors. Through one-hop inference on atomic formulas, LMPNN performs bidirectional logical message passing on each edge in the query graph. However, such a mechanism ignores the difference between the constant node with a pre-trained embedding and the variable node whose embedding needs update. Our experimental results show that passing messages to the constant node and updating its embedding is unnecessary. In addition, at the node level, LMPNN aggregates the received logical messages and updates the node embeddings in a manner similar to GIN (Xu et al., 2019). Despite the expressive power of GIN in graph structure learning, it cannot dynamically measure the importance of different messages. Whether such a GIN-like approach can capture the implicit logical dependencies between messages received by a node and between messages and that node remains unclear. By contrast, our work only performs conditional message passing on the query graph, which avoids unnecessary computation costs. Furthermore, by using the transformer (Vaswani et al., 2017) to aggregate the message and update the corresponding node embedding, our work can assign adaptive weights to elements in an input set consisting of received messages and the corresponding node and explicitly model logical dependencies between various elements.
@@
In addition to neural CQA models, some recent studies integrate symbolic information into neural models, which can be called symbolic integration models, also known as neural-symbolic models (Arakelyan et al., 2020; Xu et al., 2022; Zhu et al., 2022; Arakelyan et al., 2023; Bai et al., 2023). These models estimate the probability of whether each entity is the answer at each intermediate step in the process of multi-hop logical reasoning. This makes the size of the intermediate states for symbolic reasoning in symbolic integration models scale linearly with the size of the KG. Therefore, compared with neural CQA models where the intermediate embeddings are of fixed dimension, symbolic integration models require more computing resources and are more likely to suffer from scalability problems on large-scale KGs (Ren et al., 2023; Wang et al., 2023b).
@@ 3. Background
Since disjunctive queries can be solved in a scalable manner by transforming queries into the disjunctive normal form, it is only necessary to define query graphs for conjunctive queries. We follow previous works (Wang et al., 2023b; Ren* et al., 2020) and represent the conjunctive query as the query graph where the terms are represented as nodes connected by the atomic formulas. That is, each edge in the query graph is an atomic formula. According to the definition in Equation 2, an atomic formula contains relations, negation information, and terms. Therefore, each node in the query graph is either a constant entity or a free or existential variable, as illustrated in Figure 1.
@@
As shown in Figure 1, each edge in a query graph is an atomic formula containing the information of relation, logical negation, and terms. Passing a logical message on an edge is essentially an operation of utilizing this information to perform one-hop inference on an atomic formula. On each edge, when a node is at the head position, its neighbor is at the tail position, and vice versa. A logical message passed from a neighbor to a node is essentially an intermediate embedding of that node. Prior work (Wang et al., 2023b) proposes to obtain the intermediate embeddings of nodes by one-hop inference that maximizes the continuous truth value of the (negated) atomic formulas. Specifically, a logical message encoding function $\rho$ is proposed to perform one-hop inference. Such function has four input parameters, including neighbor embedding, relation embedding, direction information ($h\to t$ or $t\to h$), and logical negation information (00 for no negation and $1$ for with negation), and it has four cases depending on the input parameters. Given the tail embedding $t$ and relation embedding $r$ on a non-negated atomic formula, $\rho$ is formulated in the form of continuous truth value maximization to infer the head embedding $\hat{h}$:
@@
For Equations 1013, the $\lambda$ is a hyperparameter. In our work, we follow prior work (Wang et al., 2023b) and use the above logical message encoding function $\rho$ to compute messages passed to the variable nodes.
Figure 2. An illustration of the two-stage procedures of CLMPT. (a) Passing the logical message to the variable node in the neighborhood. (b) Updating the existential variable node embedding with the received messages and a transformer encoder. Similarly, such a node embedding update scheme is applied to other variable nodes, such as the free variable node $y$.
@@ 4.2. Node Embedding Conditional Update Scheme
where $Mean$ is a mean pooling layer, $TE$ is a standard transformer encoder. Through the self-attention mechanism in the transformer, we can explicitly model the complex logical dependencies between various elements in the input set and dynamically measure the importance of different elements. Figure 2 (b) shows how to obtain the updated embedding of the existential variable node $x$ through the encoding process defined by Equation 15.

### v2

@@ Abstract.
Complex Query Answering (CQA) over Knowledge Graphs (KGs) is a challenging task. Given that KGs are usually incomplete, neural models are proposed to solve CQA by performing multi-hop logical reasoning. However, most of them cannot perform well on both one-hop and multi-hop queries simultaneously. Recent work proposes a logical message passing mechanism based on the pre-trained neural link predictors. While effective on both one-hop and multi-hop queries, it ignores the difference between the constant and variable nodes in a query graph. In addition, during the node embedding update stage, this mechanism cannot dynamically measure the importance of different messages, and whether it can capture the implicit logical dependencies related to a node and received messages remains unclear. In this paper, we propose Conditional Logical Message Passing Transformer (CLMPT), which considers the difference between constants and variables in the case of using pre-trained neural link predictors and performs message passing conditionally on the node type. We empirically verified that this approach can reduce computational costs without affecting performance. Furthermore, CLMPT uses the transformer to aggregate received messages and update the corresponding node embedding. Through the self-attention mechanism, CLMPT can assign adaptive weights to elements in an input set consisting of received messages and the corresponding node and explicitly model logical dependencies between various elements. Experimental results show that CLMPT is a new state-of-the-art neural CQA model. https://github.com/qianlima-lab/CLMPT.
Knowledge Graph; Logical Reasoning; Complex Query Answering; Graph Neural Network
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 25–29, 2024; Barcelona, Spain.booktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spainisbn: 979-8-4007-0490-1/24/08doi: 10.1145/3637528.3671869ccs: Computing methodologies Reasoning about belief and knowledgeccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
Recently, neural models (Ren and Leskovec, 2020; Amayuelas et al., 2022; Zhang et al., 2021; Bai et al., 2023a) have been proposed to solve Complex Query Answering (CQA) over incomplete KGs. The complex queries that these models aim to address belong to an important subset of the first-order queries, specifically Existentially quantified First Order queries with a single free variable (EFO-1) (Wang et al., 2022, 2023b). Such EFO-1 queries contain logical operators, including conjunction, disjunction, and negation, as well as the existential quantifier ($\exists$) (Ren and Leskovec, 2020). As shown in Figure 1, one can get the corresponding EFO-1 formula and query graph given a question. The query graph is a graphical representation of the EFO-1 query, where each edge is an atomic formula that contains a predicate with a (possible) negation operator, and each node represents an input constant entity or a variable. Most neural CQA models convert the query graph into the computation graph, where logical operators are replaced with corresponding set operators. These models embed the entity set into specific vector spaces (Ren and Leskovec, 2020; Zhang et al., 2021; Chen et al., 2022; Yang et al., 2022) and execute set operations parameterized by the neural networks according to the computational graph. However, these models cannot perform well on both one-hop and multi-hop queries simultaneously. They tend to be less effective than classical neural link predictors (Trouillon et al., 2016) on one-hop atomic queries (Arakelyan et al., 2020; Wang et al., 2023b).
@@
To this end, Logical Message Passing Neural Network (LMPNN) (Wang et al., 2023b) proposes a message passing framework based on pre-trained neural link predictors. Specifically, for each edge in the query graph, LMPNN uses the pre-trained neural link predictor to infer an intermediate embedding for a node given neighborhood information. The intermediate embedding can be interpreted as a logical message passed by the neighbor on the edge. Figure 1 demonstrates the logical message passing with red/blue arrows. Then, LMPNN aggregates messages and updates node embeddings in a manner similar to Graph Isomorphism Network (GIN) (Xu et al., 2019). While effective on both one-hop and multi-hop queries, LMPNN ignores the difference between constant and variable nodes. Logical message passing can be viewed as learning an embedding for the variable by utilizing information of constant entities. It makes this embedding similar or close to the embedding of the solution entity in the embedding space of the pre-trained neural link predictor. However, as shown by the red arrows in Figure 1, LMPNN also passes messages to constant entitiy nodes with pre-trained embeddings and updates their embeddings, which we believe is unnecessary and experimentally demonstrated. In addition, for node embedding updating, despite the expressive power of GIN in graph structure learning (Xu et al., 2019), it cannot dynamically measure the importance of different logical messages. Furthermore, it is unclear whether the GIN-like approach can capture the implicit complex logical dependencies between messages received by a node and between messages and that node. Explicitly modeling the above implicit logical dependencies while being able to measure the importance of different messages dynamically is an open challenge we take on in this paper.
@@
We propose Conditional Logical Message Passing Transformer (CLMPT), a special Graph Neural Network (GNN) for CQA. Each CLMPT layer has two stages: (1) passing the logical message of each node to all of its neighbors that are not constant entity nodes; (2) updating the embedding of each variable node based on the logical messages received by the variable node. This means that during forward message passing, CLMPT does not use the pre-trained neural link predictor to infer the intermediate embeddings (i.e. logical messages) for constant entity nodes with pre-trained embeddings and does not update constant node embeddings. We call such a mechanism Conditional Logical Message Passing, where the condition depends on whether the node type is constant or variable. We empirically verified that it can reduce computational costs without affecting performance. Furthermore, CLMPT uses the transformer (Vaswani et al., 2017) to aggregate messages and update the corresponding node embedding. The self-attention mechanism in the transformer can dynamically measure the importance of different elements in the input and capture the interactions between any two elements. It allows us to explicitly model the logical dependencies between messages received by a node and between messages and that node. We conducted experiments on three popular KGs: FB15k (Bordes et al., 2013), FB15k-237 (Toutanova and Chen, 2015), and NELL995 (Xiong et al., 2017a). The experimental results show that CLMPT can achieve a strong performance.
The main contributions of this paper are summarized as follows:
  • We propose a conditional logical message passing mechanism, which can effectively reduce the computational costs and even improve the model performance to some extent.
  • We propose a transformer-based node embedding update scheme. It can dynamically measure the importance of different intermediate embeddings of a variable node and capture possible implicit dependencies between these embeddings.
  • Extensive experiments results on several benchmark datasets show that CLMPT is a new state-of-the-art neural CQA model.
@@ 2. Related Work
Complex queries over KGs can be regarded as one-hop atomic queries combined with existential first-order logic operators. The scope of complex queries that existing works can answer is expanded from conjunctive queries (Hamilton et al., 2018; Kotnis et al., 2021), to Existential Positive First-Order (EPFO) queries (Ren* et al., 2020; Arakelyan et al., 2020), and more recently to the existential first-order (Ren and Leskovec, 2020), or more specifically to EFO-1 (Wang et al., 2022, 2023b) and $\text{EFO}_{k}$ queries (Yin et al., 2023). Recently, neural query embedding (QE) models have been proposed to answer complex queries. Most QE models embed entity sets into specific continuous vector spaces with various forms, including geometric shapes (Choudhary et al., 2021b; Amayuelas et al., 2022; Hamilton et al., 2018; Ren* et al., 2020; Zhang et al., 2021; Liu et al., 2021; Nguyen et al., 2023; Bai et al., 2022), probabilistic distributions (Ren and Leskovec, 2020; Yang et al., 2022; Choudhary et al., 2021a), fuzzy logic (Chen et al., 2022), and bounded histograms on uniform grids (Wang et al., 2023a). They represent the complex query in the form of the computation graph and perform neural set operations in the corresponding embedding spaces according to the computation graph. Such neural set operations are usually modeled by complex neural networks (Tolstikhin et al., 2021; Vaswani et al., 2017). In spite of getting query embedding with neural set operations, other QE models represent the queries as graphs to embed queries with GNNs (Daza and Cochez, 2020; Alivanistos et al., 2022) or Transformers (Kotnis et al., 2021; Liu et al., 2022; Wang et al., 2023c; Bai et al., 2023c). However, most of these graph representation learning models only focus on EPFO queries, and how they handle the logical negation operator is unclear. Moreover, the aforementioned neural models are still less effective than a simple neural link predictor (Trouillon et al., 2016) on one-hop atomic queries.
@@
There is also a class of neural CQA models based on pre-trained neural link predictors. CQD-CO (Arakelyan et al., 2020) proposes a continuous optimization strategy that uses pre-trained neural link predictors (Trouillon et al., 2016) and logical t-norms (Klement et al., 2013) to solve complex queries. However, according to (Wang et al., 2023b), CQD-CO performs badly on negative queries. Q2T (Xu et al., 2023) is based on graph transformers (Ying et al., 2021) and pre-trained neural link predictors to encode complex queries. LMPNN (Wang et al., 2023b), which is most related to our work, proposes a logical message passing framework based on pre-trained neural link predictors. Through one-hop inference on atomic formulas, LMPNN performs bidirectional logical message passing on each edge in the query graph. However, such a mechanism ignores the difference between the constant and variable nodes. The essence of logical message passing lies in leveraging the pre-trained entity embeddings corresponding to constant nodes to infer variable node embeddings. It aligns the variable node embeddings more closely to the embeddings of solution entities in the embedding space of the pre-trained neural link predictor. It is important to note that using the intermediate states of the variable nodes to update the embeddings of constant nodes can introduce noise. This occurs because constant nodes possess stable, pre-trained embeddings corresponding to specific entities, which do not require updates via message passing. Therefore, updating constant node embeddings with information from variable nodes disrupts the integrity of these established embeddings. In addition, at the node level, LMPNN aggregates the received logical messages and updates the node embeddings in a manner similar to GIN (Xu et al., 2019). Despite the expressive power of GIN in graph structure learning, it cannot dynamically measure the importance of different messages. Whether such a GIN-like approach can capture the implicit logical dependencies between messages received by a node and between messages and that node remains unclear. By contrast, our work only performs conditional message passing on the query graph, which avoids unnecessary computation costs. Furthermore, by using the transformer (Vaswani et al., 2017) to aggregate the message and update the corresponding node embedding, our work can assign adaptive weights to elements in an input set consisting of received messages and the corresponding node and explicitly model logical dependencies between various elements.
@@
In addition to neural CQA models, some recent studies integrate symbolic information into neural models, which can be called symbolic integration models, also known as neural-symbolic models (Arakelyan et al., 2020; Xu et al., 2022; Zhu et al., 2022; Arakelyan et al., 2023; Bai et al., 2023b). These models estimate the probability of whether each entity is the answer at each intermediate step in the process of multi-hop logical reasoning. This makes the size of the intermediate states for symbolic reasoning in symbolic integration models scale linearly with the size of the KG. Therefore, compared with neural CQA models where the intermediate embeddings are of fixed dimension, symbolic integration models require more computing resources and are more likely to suffer from scalability problems on large-scale KGs (Ren et al., 2023; Wang et al., 2023b).
@@ 3. Background
Since disjunctive queries can be solved in a scalable manner by transforming queries into the disjunctive normal form, it is only necessary to define query graphs for conjunctive queries. We follow previous works (Wang et al., 2023b; Ren* et al., 2020) and represent the conjunctive query as the query graph where the terms are represented as nodes connected by the atomic formulas. That is, each edge in the query graph is an atomic formula. According to the definition in Equation 2, an atomic formula contains relations, negation information, and terms. Therefore, each node in the query graph is either a constant entity or a free or existential variable, as illustrated in Figure 1. In this paper, the term ”constants” has specific meanings depending on the context. When discussing an EFO-1 query, ”constants” refers to the input constant entities within that query. In the context of the query graph, ”constants” denotes the constant nodes that correspond to these input constant entities.
@@
As shown in Figure 1, each edge in a query graph is an atomic formula containing the information of relation, logical negation, and terms. Passing a logical message on an edge is essentially an operation of utilizing this information to perform one-hop inference on an atomic formula. On each edge, when a node is at the head position, its neighbor is at the tail position, and vice versa. A logical message passed from a neighbor to a node is essentially an intermediate embedding of that node. Prior work (Wang et al., 2023b) proposes to obtain the intermediate embeddings of nodes by one-hop inference that maximizes the continuous truth value of the (negated) atomic formulas. Specifically, a logical message encoding function $\rho$ is proposed to perform one-hop inference. Such a function has four input parameters, including neighbor embedding, relation embedding, direction information ($h\to t$ or $t\to h$), and logical negation information (00 for no negation and $1$ for with negation), and it has four cases depending on the input parameters. Given the tail embedding $t$ and relation embedding $r$ on a non-negated atomic formula, $\rho$ is formulated in the form of continuous truth value maximization to infer the head embedding $\hat{h}$:
@@
For Equations 1013, $\lambda$ is a hyperparameter. In our work, we follow prior work (Wang et al., 2023b) and use the above logical message encoding function $\rho$ to compute messages passed to the variable nodes.
Figure 2. An illustration of the two-stage procedures of CLMPT at the $l$-th layer. (a) Passing the logical message to the variable node in the neighborhood. (b) Updating the existential variable node embedding with the received messages and a transformer encoder. Similarly, such a node embedding update scheme is applied to other variable nodes, such as the free variable node $y$.
@@ 4.2. Node Embedding Conditional Update Scheme
where $Mean$ is a mean pooling layer. $TE$ is a standard transformer encoder stacked with multiple layers of transformer encoder blocks, where the attention computation approach is consistent with the bidirectional multi-head self-attention of the vanilla transformer (Vaswani et al., 2017). Specifically, given an input embedding set $X$, it is first projected to query ($Q$), key ($K$), and value ($V$) matrices through a linear projection such that $Q=XW^{Q}$, $K=XW^{K}$ and $V=XW^{V}$ respectively. Then, the self-attention can be compute via
(16) $\begin{split}Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{K}}})V,\end{split}$
where $d_{K}$ represents the dimension of $K$, and $W^{Q}$, $W^{K}$, $W^{V}$ are parameter matrices. We use multi-head attention, which concatenates multiple instances of Equation 16 and then obtains the output through a linear projection. It is worth noting that since the input to Equation 15 is a set consisting of the node embedding and the messages rather than a sequence, we do not use any positional encoding. This means that our architecture is permutation invariant. Through the self-attention mechanism, we can explicitly model the complex logical dependencies between various elements in the input set and dynamically measure the importance of different elements. Figure 2 (b) shows how to obtain the updated embedding of the existential variable node $x$ through the encoding process defined by Equation 15.